# Simulating the shaping of point-symmetric structures in the jittering jets explosion mechanism

### v1

@@ Abstract
We conduct three-dimensional hydrodynamical simulations of core-collapse supernovae by launching several pairs of jets into a collapsing core model, and show that the jittering jets explosion mechanism (JJEM) can form a point-symmetric morphology that accounts for observed morphologies of many core-collapse supernovae (CCSN) remnants. Point-symmetric morphologies are composed of pairs of opposite structures around the center of the CCSN remnant. In the JJEM, the newly born neutron star launches several to a few tens of pairs of jets with stochastically varying directions, and these jets explode the star. In the simulations with the FLASH numerical code, we launch pairs of jets with varying directions, energies, opening angles, and durations into the massive stellar core and follow their evolution for about two seconds. We show that the jets form pairs of opposite filaments, clumps, bubbles, and lobes, namely, prominent point-symmetric morphologies. The interaction of the jets with the core leads to vigorous Rayleigh-Taylor instabilities and excites many vortices, which also shape clumps and filaments. Our results strengthen the claim that the JJEM is the primary explosion mechanism of CCSNe.
@@ 2.3 Simulated cases
$\theta_{\rm j}\ [\circ]$ $50$ $50$ $50$
ϕj[]subscriptitalic-$\phi_{\rm j}\ [\circ]$ 00 $120$ $240$
$\alpha_{\rm j}\ [\circ]$ $10$ $10$ $10$
@@
$\theta_{\rm j}\ [\circ]$ $50$ $50$ $50$ 00 00
ϕj[]subscriptitalic-$\phi_{\rm j}\ [\circ]$ 00 $120$ $240$ 00 $20$
$\alpha_{\rm j}\ [\circ]$ $10$ $10$ $10$ $10$ $20$
@@ Appendix A Initial parameters of simulations
$\theta_{\rm j}\ [\circ]$ 00 $40$ $80$
ϕj[]subscriptitalic-$\phi_{\rm j}\ [\circ]$ 00 00 00
$\alpha_{\rm j}\ [\circ]$ $10$ $10$ $10$
@@
$\theta_{\rm j}\ [\circ]$ 00 $40$ $80$
ϕj[]subscriptitalic-$\phi_{\rm j}\ [\circ]$ 00 00 00
$\alpha_{\rm j}\ [\circ]$ $40$ $10$ $10$
@@
$\theta_{\rm j}\ [\circ]$ $50$ $50$ $50$ 00 00
ϕj[]subscriptitalic-$\phi_{\rm j}\ [\circ]$ 00 $120$ $240$ 00 $20$
$\alpha_{\rm j}\ [\circ]$ $10$ $10$ $10$ $10$ $5$
@@
$t_{\rm j}({\rm start})$ $0.00$ $0.07$ $0.14$ $0.21$ $0.28$ $0.35$ $0.42$
$[\rm s]$
$t_{\rm j}{\rm(end)}$ $0.05$ $0.12$ $0.19$ $0.26$ $0.33$ $0.40$ $0.47$
$[\rm s]$
$\theta_{\rm j}$ $50$ $50$ $50$ 00 00 $50$ $30$
[]delimited-[][$\circ$][ ∘ ]
$\phi_{\rm j}$ 00 $120$ $240$ 00 $20$ $240$ $70$
[]delimited-[][$\circ$][ ∘ ]
$\alpha_{\rm j}$ $10$ $10$ $10$ $10$ $20$ $10$ $10$
[]delimited-[][$\circ$][ ∘ ]
$\dot{m}_{\rm j}$ $1.8$ $1.8$ $1.8$ $1.8$ $1.8$ $1.8$ $1.2$
$[10^{32}{~{}\rm g}{~{}\rm s}^{-1}]$
$E_{\rm k,2j}$ $2.25$ $2.25$ $2.25$ $2.25$ $2.25$ $2.25$ $1.5$
$[10^{50}\ \rm erg]$

### v2

@@ Abstract
We conduct three-dimensional hydrodynamical simulations of core-collapse supernovae by launching several pairs of jets into a collapsing core model and show that the jittering jets explosion mechanism (JJEM) can form a point-symmetric morphology that accounts for observed morphologies of many core-collapse supernovae (CCSN) remnants. Point-symmetric morphologies are composed of pairs of opposite structures around the center of the CCSN remnant. In the JJEM, the newly born neutron star launches several to a few tens of pairs of jets with stochastically varying directions, and these jets explode the star. In the simulations with the FLASH numerical code, we launch pairs of jets with varying directions, energies, opening angles, and durations into the massive stellar core and follow their evolution for about two seconds. We show that the jets form pairs of opposite filaments, clumps, bubbles, and lobes, namely, prominent point-symmetric morphologies. The interaction of the jets with the core leads to vigorous Rayleigh-Taylor instabilities and excites many vortices, which also shape clumps and filaments. Our results strengthen the claim that the JJEM is the primary explosion mechanism of CCSNe.
@@ 2.3 Simulated cases
$\theta_{\rm j}\ [\rm deg]$ $50$ $50$ $50$
$\phi_{\rm j}\ [\rm deg]$ 00 $120$ $240$
$\alpha_{\rm j}\ [\rm deg]$ $10$ $10$ $10$
@@
$\theta_{\rm j}\ [\rm deg]$ $50$ $50$ $50$ 00 00
$\phi_{\rm j}\ [\rm deg]$ 00 $120$ $240$ 00 $20$
$\alpha_{\rm j}\ [\rm deg]$ $10$ $10$ $10$ $10$ $20$
@@ Appendix A Initial parameters of simulations
$\theta_{\rm j}\ [\rm deg]$ 00 $40$ $80$
$\phi_{\rm j}\ [\rm deg]$ 00 00 00
$\alpha_{\rm j}\ [\rm deg]$ $10$ $10$ $10$
@@
$\theta_{\rm j}\ [\rm deg]$ 00 $40$ $80$
$\phi_{\rm j}\ [\rm deg]$ 00 00 00
$\alpha_{\rm j}\ [\rm deg]$ $40$ $10$ $10$
@@
$\theta_{\rm j}\ [\rm deg]$ $50$ $50$ $50$ 00 00
$\phi_{\rm j}\ [\rm deg]$ 00 $120$ $240$ 00 $20$
$\alpha_{\rm j}\ [\rm deg]$ $10$ $10$ $10$ $10$ $5$
@@
$t_{\rm j}({\rm start})[\rm s]$ $0.00$ $0.07$ $0.14$ $0.21$ $0.28$ $0.35$ $0.42$
$t_{\rm j}{\rm(end)}[\rm s]$ $0.05$ $0.12$ $0.19$ $0.26$ $0.33$ $0.40$ $0.47$
$\theta_{\rm j}[\rm deg]$ $50$ $50$ $50$ 00 00 $50$ $30$
$\phi_{\rm j}[\rm deg]$ 00 $120$ $240$ 00 $20$ $240$ $70$
$\alpha_{\rm j}[\rm deg]$ $10$ $10$ $10$ $10$ $20$ $10$ $10$
$\dot{m}_{\rm j}[10^{32}{~{}\rm g}{~{}\rm s}^{-1}]$ $1.8$ $1.8$ $1.8$ $1.8$ $1.8$ $1.8$ $1.2$
$E_{\rm k,2j}[10^{50}\ \rm erg]$ $2.25$ $2.25$ $2.25$ $2.25$ $2.25$ $2.25$ $1.5$